# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 4.5 Heterogeneity analysis
The identification result in equation (4.2) can be altered slightly to show that $ATT(2)$ is an aggregation of covariate-specific $2\times 2$ DiD estimands:
@@
This not only demonstrates the building block structure of conditional designs, it also connects the strategies we discussed for estimating $ATT(2)$ itself to the underlying treatment effect heterogeneity by covariate values. When this heterogeneity is of interest in its own right it can also be targeted, identified, and estimated (at least under some additional conditions).
When all covariates are discrete, estimating $ATT_{X_{i}}(2)$ parameters is fairly straightforward. One can form saturated partitions $x_{k},\leavevmode\nobreak\ k=1,\dots,K$, and then subset the data only to contain information from the units from the specified partition $k$. At this stage, the analysis is analogous to the unconditional DiD setup (except that you need to repeat this step from all partitions of interest), as each $ATT_{x_{k}}(2)$ is identified by a (conditional on a discrete-variable) comparison of means, i.e.,
@@
When some covariates are continuous, or there are arguably too many partitions that make each of them contain few observations, one can still identify the $ATT_{X_{i}}(2)$’s, though they are harder to estimate without auxiliary assumptions. In such cases, it is customary to identify and estimate a more aggregated conditional ATT parameter than $ATT_{X_{i}}(2)$. For instance, one may want to assess if the effect of Medicaid expansion on adult mortality rates is higher (or lower) in counties with an unemployment rate above the median than those below. Similar partitions could be made for any other covariates. To formalize this notion of “partition specific” ATT, let $\text{PART}(X_{i})$ be some user-specified partition of the covariate space such that $\text{PART}(X_{i})\in\{1,2,\dots,K\}$, and define
@@
Thus, for heterogeneity analysis with covariates, one can partition the data with a user-specified partition map $\text{PART}(X_{i})$ and then, within each of these partitions, use similar arguments as we used to establish (4.2) to guarantee that each partition-specific ATT is identified (which is precisely what we did to get $ATT_{k}(2)$ above). Regarding estimation and inference, one can use regression adjustment, inverse probability weighting, or doubly robust estimators as in Section 4.4; the difference is that you need to “localize” the analysis for each partition.
@@ 5 DiD designs with multiple time periods
The previous sections focused on fundamental identification issues and estimation approaches for $2\times 2$ building blocks, but generally did not build them into anything because they targeted the only feasible $ATT(t)$, $ATT(2)$. DiD designs with multiple periods are notably different. They can target $ATT$s in each period after treatment to trace out dynamics, and they can produce cross-group outcome trends from before treatment starts to evaluate the plausibility of parallel trends. Multiple periods also admit the possibility of more complex treatment variation. We will focus on staggered-timing designs in which treatment “turns on” at different times for different units and stays on, as the Medicaid expansion did. It is possible, however, to use DiD methods to study treatments that “turn off” or occur more than once, though some modifications are warranted; see, e.g., deChaisemartin2020; deChaisemartin2023b-intertemporal-treatments.
We need to expand the notation to define the relevant concepts with multiple periods. When treatment can only happen at one point in time (so we still have only two treatment groups with $D_{i}$ equal to one or zero), the only changes we need to make are to acknowledge that time runs from $t=1,2,\dots,T$, and that the map between potential outcomes and observed outcomes is
@@ 5.1 Simple event studies $\mathbf{(2\times T)}$
The term “event study” refers to estimating and reporting effects across a range of time periods before and after treatment. A design with one treatment timing group and multiple time periods ($2\times T$) is the simplest case in which to discuss event studies. We thus expand our analysis of the 2014 Medicaid expansion group to include data from 2009 to 2019. We report population-weighted results for brevity, as this is arguably more policy-relevant in the Medicaid expansion context. Figure 2 plots the time series of the weighted mortality rates for the 2014 and post-2019 expansion counties. It is the analog of Table 2 in the sense that it presents the raw data elements necessary to construct event study estimates. The treatment year 2014 naturally divides the x-axis into two windows: post-treatment (2014-2019) and pre-treatment (2009-2013). An event study constructs DiD-type estimates in both windows, but they have different interpretations.
@@
Notes: This figure shows county population-weighted average mortality rates for adults ages 20-64 by expansion category from 2009 to 2019.
@@ 5.1.1 Event study estimates in the post-treatment periods
The average change of $Y_{i,t}(0)$ from $Y_{i,t=g-1}(0)$ is the same between treated and comparison units for all post-treatment periods $t\geq g$, i.e.,
@@
Assumption PT-ES suggests that learning about long-run effects requires stronger assumptions than learning about short-run effects. That is, to identify the average effect of Medicaid expansion in 2019 among expansion counties, Assumption PT-ES requires parallel trends to hold in every year from 2014 to 2019. On the other hand, if we are interested in learning only about short-run effects, say effects up until 2015, we would require it to hold from 2014 and 2015 only.
If Assumption PT-ES holds (as well as no-anticipation), then each $ATT(t)$ is identified by a DiD comparison between period $g-1$ and $t$ exactly like in (3.5), and they can be estimated by the familiar comparison of four sample averages:
@@
Figure 3 plots (weighted) event study estimates for the 2014 Medicaid expansion group. $\widehat{ATT}(t)$’s lie to the right of the vertical dashed line. The estimate for event-time 0, which means that $t=g=2014$ in this case equals the $2\times 2$ results from table 2 ($-2.7$). The other estimates have the same DiD form: comparisons of cross-group changes in different post-2014 years ($t$) but always relative to 2013 ($g-1$). The point estimates do not suggest large mortality effects from Medicaid expansion among expansion counties.
@@
Notes: This figure shows the event study estimates in the $2\times T$ case, comparing counties that expanded in 2014 to counties that have not yet expanded by 2019, using the regression adjustment estimator from Callaway2021. The outcome variable is the crude mortality rate for adults ages 20-64, and the standard errors are clustered at the county level. The point estimate is reported by the circles, and both point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines.
@@ 5.1.2 Event study estimates in the pre-periods
The $\tau_{-k}$ terms are usually called “differential trends” or “pre-trends”, and they appear to the left of the vertical dashed line in Figure 3. The credibility of event study analyses rests in large part on finding small estimates of $\tau_{-k}$, but Figure 3 illustrates how challenging it can be to draw conclusions from informally looking at these pre-trends. No individual $\tau_{-k}$ is statistically significant, so we fail to reject the null hypothesis that the pre-trend estimates equal zero (individually or jointly). This kind of result is often interpreted to mean that parallel trends is plausible. But the $\tau_{-k}$’s also tend to be positive with a mean of about 2.3, which is larger in magnitude than all post-period point estimates but one. Sometimes, this kind of result is used to argue that parallel trends may not be plausible. Ultimately, do these results support or refute the plausibility of parallel trends?
@@
Recent methodological work suggests three practical lessons about how to handle pre-trend estimates in terms of assessing the plausibility of parallel trends; see, e.g., Bilinski2018, Manski2018, KahnLang2019, Roth2022, Rambachan2023, Dette_Schumann_JBES_2024, and Freyaldenhoven_etal_2024. The most fundamental, which will tend to shape language more than practice, is that Assumption PT-ES is not testable, as it only makes restrictions on untreated potential outcomes in post-treatment periods, $t\geq g$. Under no-anticipation, pre-trends measure something with the same form as parallel trends—differences in untreated outcome trends between treated and untreated units—but they necessarily measure it in the “wrong” periods $t<g$. This does not mean parallel pre-trends are not informative; it just means they are not the same as the parallel trends assumption in Assumption PT-ES. If the factors that drive pre-trends also matter for post-treatment trends in $Y_{i,t}(0)$, then pre-treatment coefficients close to zero support the plausibility of Assumption PT-ES. In such cases, large pre-trends may cast doubt on it.
It is important to note that pre-trends are not necessarily informative about parallel trends in post-treatment periods. Sometimes, looking too far before treatment starts is not informative about parallel trends in post-treatment periods, as the economic environment may be too different from time periods near the treatment date. Ultimately, how informative pre-trends are for the plausibility of Assumption PT-ES is case-specific, though we would prefer to have parallel pre-trends. In fact, when it is plausible to strengthen Assumption PT-ES to hold in all periods (and not only in post-treatment periods), one can construct estimators for the $ATT(t)$ that are more precise than those in (5.2); see, e.g., Borusyak2023, Gardner2021, Harmon2024, Marcus_Santanna_2021, Wooldridge2021, and Chen_SantAnna_Xie_effientDiD_2024. In fact, Chen_SantAnna_Xie_effientDiD_2024 derive the semiparametric efficiency bound and propose simple-to-use plug-in estimators with the shortest possible confidence intervals one can get without introducing additional modeling restrictions. Such estimators weight pre-treatment in a non-uniform manner and explore the correlation structure of changes in outcomes across periods and between groups to explore all the empirical content of the identification assumptions efficiently. In such cases, though, one would require parallel trends to hold pre-treatment periods, parallel pre-trends becomes testable (as the model is over-identified), and $\tau_{-k}$ could be used to assess its plausibility directly; see Marcus_Santanna_2021 and Chen_SantAnna_Xie_effientDiD_2024 for a discussion.
@@
The second lesson is that statistical precision shapes the usefulness of pre-trend estimates. The hypothesis tests for parallel pre-trends in figure 3 are low-powered to detect practically important violations (Roth2022, and Freyaldenhoven_etal_2024). The $\tau_{-k}$ estimates fail to rule out flat pre-trends, but as we shall see below, they also fail to rule out large pre-trends that would indicate serious bias in the $\widehat{ATT}(t)$s. They simply do not say very much. [^16] Another possibility is that very precisely estimated pre-trends are distinguishable from zero yet also rule out even small violations. The magnitude of the violations also matters: a precisely estimated but small violation of parallel trends in pre-treatment periods is “better” than an imprecisely potentially large estimated pre-trend that does not rule out zero. Roth2022 discusses how conditioning the analysis on these kinds of low-powered tests can exacerbate biases and should be interpreted with care.
@@
The third lesson is that researchers can make better use of pre-trend estimates by taking a stand on the size of plausible and/or problematic parallel trends violations. For example, Bilinski2018 propose selecting a value for differential trends that would fully explain the estimated treatment effects and then using it as the null hypothesis instead of zero when conducting statistical tests of the estimated pre-trends; see also Dette_Schumann_JBES_2024. Rambachan2023 develop inference methods for an approach that bounds the $ATT(t)$s under assumptions about the maximum size of parallel trends violations based on pre-treatment periods (see also Manski2018). One can either select a magnitude using contextual knowledge or set it equal to a multiple of the largest period-to-period pre-trend estimate. One then constructs an identified set that contains $ATT(t)$ not under parallel trends but under the weaker assumption that parallel trends is not violated by more than this pre-determined maximum in each post-period. Importantly, Rambachan2023 also show how to use the estimated covariance matrix of the event study estimates to construct confidence intervals around the set. [^17] In cases where a measured variable accounts for the pre-trends, Freyaldenhoven2019 develop two-stage-least-squares estimators that recover $ATT(t)$ parameters by extrapolating the pre-period relationship between outcomes and the covariate into the post-period. Given the existence of these methods, which are theoretically grounded in how to consider violations of pre-trends, we caution producers (and consumers) of DiD work against the common practice of only using a simple “eye-test” for whether pre-trends differ substantially from zero.
@@
In our application, Rambachan2023’s method underscores how little information the pre-trend estimates convey. The largest one-period pre-trend is between event-time -5 and -4, when outcomes fall by roughly four deaths more in the expansion group versus the non-expansion group. If we assume that parallel trend violations are no bigger than this, the identified set for $ATT(2014)$ is $-2.6\pm 4=[-6.6,1.4]$, and given the size of the pre-period standard errors, we obtain a robust confidence interval of $[-11.1,5.1]$, and interval that spans implausibly large effects in both directions. Assuming smaller parallel trend violations would shrink this interval, but applying the method to subsequent event-times widens it. In general, Rambachan2023 provide a flexible method to use information about potential parallel trends violations drawn either from external knowledge or the $\tau_{-k}$ estimates, as well as the precision of the pre-trend estimates, to gauge (statistically) the robustness of the $\widehat{ATT}(t)$’s.
@@
When pre-trends suggest that Assumption PT-ES is not plausible, a way forward is to assume that it only holds after conditioning on covariates and proceeding similarly to what we described in Section 4.4; we discuss this path in detail in Section 5.1.4. Alternatively, one can attempt to parametrically model the violations of parallel trends, usually by including unit-specific linear trends; see, e.g., Mora_Reggio_ER_2019, Wooldridge2021, Lee_Wooldridge_2023, and Freyaldenhoven_etal_2024. We note, however, that the practice of using unit-specific linear trends deviates from the standard DiD procedures: it relies on alternative identification assumptions involving an explicit parametric model for unit-specific trends. We also note that sensitivity analysis procedures that do not rely on such models are available, e.g., Rambachan2023, and we encourage practitioners to consider them.

### v2

@@ 4.5 Heterogeneity analysis
A different motivation for using covariates is to estimate heterogeneous treatment effects by the values of $X_{it}$. The identification result in equation (4.2) can be altered slightly to show that $ATT(2)$ is simply an aggregation of covariate-specific $2\times 2$ DiD estimands:
@@
This not only demonstrates the building block structure of conditional designs, but also connects the strategies we discussed for estimating $ATT(2)$ itself to the underlying treatment effect heterogeneity by covariate values. When this heterogeneity is of interest in its own right it can also be targeted, identified, and estimated (at least under some additional conditions).
When all covariates are discrete, estimating $ATT_{X_{i}}(2)$ parameters is fairly straightforward. One can form saturated partitions $x_{k},\leavevmode\nobreak\ k=1,\dots,K$, and then subset the data to contain only information from the units from the specified partition $k$. At this stage, the analysis is analogous to the unconditional DiD setup (except that you need to repeat this step from all partitions of interest), as each $ATT_{x_{k}}(2)$ is identified by a (conditional on a discrete-variable) comparison of means’ that is,
@@
When some covariates are continuous or there are so many partitions that each one contains few observations, one can still identify the $ATT_{X_{i}}(2)$’s, though they are hard to estimate unless auxiliary assumptions hold. In such cases, it is customary to identify and estimate a more aggregated conditional ATT parameter than $ATT_{X_{i}}(2)$. For instance, one may want to assess if the effect of Medicaid expansion on adult mortality rates is higher (or lower) in counties with an unemployment rate above the median than it is in those below. Similar partitions could be made for any other covariates. To formalize this notion of “partition specific” ATT, let $\text{PART}(X_{i})$ be some user-specified partition of the covariate space such that $\text{PART}(X_{i})\in\{1,2,\dots,K\}$, and define
@@
Thus, for heterogeneity analysis with covariates, one can partition the data with a user-specified partition map $\text{PART}(X_{i})$ and then, within each of these partitions, use arguments like the ones we used to establish (4.2) to guarantee that each partition-specific ATT is identified (which is precisely what we did to get $ATT_{k}(2)$ above). Regarding estimation and inference, one can use regression adjustment, inverse probability weighting, or doubly robust estimators as in Section 4.4; the difference is that these are implemented “locally” on each partition.
@@ 5 DiD designs with multiple time periods
The previous sections focused on fundamental identification issues and estimation approaches for $2\times 2$ building blocks, but generally did not build them into anything because they targeted the only feasible $ATT(t)$, $ATT(2)$. DiD designs with multiple periods are notably different. They can target $ATT$s in each period after treatment to trace out dynamics, and they can produce cross-group outcome trends from before treatment starts to evaluate the plausibility of parallel trends. Multiple periods also admit the possibility of more complex treatment variation. We will focus on staggered-timing designs in which treatment “turns on” at different times for different units and stays on, as the Medicaid expansion did. It is possible, however, to use DiD methods to study treatments that “turn off” or occur more than once, though some modifications are warranted; see, for example, deChaisemartin2020; deChaisemartin2023b-intertemporal-treatments.
We need to expand the notation to define the relevant concepts with multiple periods. When treatment can happen only at one point in time (so we still have only two treatment groups with $D_{i}$ equal to one or zero), the only changes we need to make are to acknowledge that time runs from $t=1,2,\dots,T$, and that the map between potential outcomes and observed outcomes is
@@ 5.1 Simple event studies $\mathbf{(2\times T)}$
The term “event study” refers to estimating and reporting effects across a range of time periods before and after treatment. A design with one treatment timing group and multiple time periods ($2\times T$) is the simplest case in which to discuss event studies. We thus expand our analysis of the 2014 Medicaid expansion group to include data from 2009 to 2019. We report population-weighted results for brevity. Figure 2 plots the time series of the weighted mortality rates for the 2014 and post-2019 expansion counties. It is the analog of Table 2 in the sense that it presents the raw data elements necessary to construct event study estimates. The treatment year 2014 naturally divides the x-axis into two windows: post-treatment (2014-2019) and pre-treatment (2009-2013). An event study constructs DiD-type estimates in both windows, but they have different interpretations.
@@
Notes: This figure shows county population-weighted average mortality rates for adults ages 20-64 for 978 counties that expanded Medicaid in 2014 and 1,222 counties that did not expand Medicaid by 2019 from 2009 to 2019.
@@ 5.1.1 Event study estimates in the post-treatment periods
The average change of $Y_{i,t}(0)$ from $Y_{i,t=g-1}(0)$ is the same between treated and comparison units for all post-treatment periods $t\geq g$; that is,
@@
Assumption PT-ES suggests that learning about long-run effects requires stronger assumptions than learning about short-run effects. That is, to identify the average effect of Medicaid expansion in 2019 among expansion counties, Assumption PT-ES requires parallel trends to hold in every year from 2014 to 2019. On the other hand, if we are interested in learning only about short-run effects—say effects up until 2015—we would require it to hold from 2014 and 2015 only.
If Assumption PT-ES holds (as well as no-anticipation), then each $ATT(t)$ is identified by a DiD comparison between period $g-1$ and $t$, as in (3.5), and the $ATT(t)$s can be estimated by the familiar comparison of four sample averages:
@@
Figure 3 plots (weighted) event study estimates for the 2014 Medicaid expansion group. $\widehat{ATT}(t)$’s lie to the right of the vertical dashed line. The estimate for event-time 0 (that is, $t=g=2014$) equals the $2\times 2$ results from Table 2 ($-2.6$). The other estimates have the same DiD form: comparisons of cross-group changes in different post-2014 years ($t$) but always relative to 2013 ($g-1$). The point estimates do not suggest large mortality effects from Medicaid expansion among expansion counties.
@@
Notes: This figure shows the population-weighted event study estimates in the $2\times T$ case, comparing 978 counties in states that expanded in 2014 to 1,222 counties in states that had not yet expanded by 2019. It uses the unconditional estimator from Callaway2021. The outcome variable is the crude mortality rate for adults ages 20-64, and the standard errors are clustered at the county level. The point estimate is reported by the circles, and both point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines (see section 5.1.3).
@@ 5.1.2 Event study estimates in the pre-periods
The $\tau_{-k}$ terms are usually called “differential trends” or “pre-trends,” and they appear to the left of the vertical dashed line in Figure 3. The credibility of event study analyses rests in large part on finding small estimates of $\tau_{-k}$, but Figure 3 illustrates how challenging it can be to draw conclusions from informally looking at these pre-trends. No individual $\tau_{-k}$ is statistically significant, so we fail to reject the null hypothesis that the pre-trend estimates equal zero (individually or jointly). This kind of result is often interpreted to mean that parallel trends holds. But the $\tau_{-k}$’s also tend to be positive with a mean of about 2.3, which is larger in magnitude than all post-period point estimates except one. Sometimes, this kind of result is used to argue that parallel trends fails. Do these results support or refute parallel trends?
@@
Recent methodological work suggests three practical lessons about using pre-trend estimates to assess parallel trends; examples include Bilinski2018, Manski2018, KahnLang2019, Roth2022, Rambachan2023, Dette_Schumann_JBES_2024, and Freyaldenhoven_etal_2024. The most fundamental, which will tend to shape language more than practice, is that Assumption PT-ES is not testable, as it only makes restrictions on untreated potential outcomes in post-treatment periods, $t\geq g$. Under no-anticipation, pre-trends do measure differences in untreated outcome trends between treated and untreated units, but they necessarily measure them in the “wrong” periods $t<g$. This does not mean parallel pre-trends are not informative; it just means they are not the same as the parallel trends assumption in Assumption PT-ES.
When specifically are observed pre-trends uninformative about the parallel trends assumptions required for identification? One case is when pre-trends are measured too far before treatment starts. The conditions or shocks that jointly shape treatment decisions and untreated outcomes may differ many periods before treatment, but this need not imply that they generate bias in the periods after treatment. A second case, discussed in Section 3.2, is when units select into treatment based on time-varying pre-treatment unobservables. As treatment selection depends on pre-treatment unobservables, non-parallel pre-trends may appear as a consequence of the selection mechanism, though parallel trends may still be plausible. Ghanem2023 show that in this case, a necessary condition for parallel trends is that the untreated potential outcomes satisfy a martingale property. When this martingale property does not hold, parallel trends cannot hold either, regardless of the shape of pre-trends. In such situations, it is worthwhile to assess the potential bias components of the DiD estimator using the “selection-aware” benchmarking tools presented in Ghanem2023, as well as to resort to their theory-based templates for justifying parallel trends based on selection mechanisms. Ultimately, how informative pre-trends are for the plausibility of Assumption PT-ES is case-specific, though we would of course always prefer to have parallel pre-trends.
When Assumption PT-ES holds in all periods (pre- and post-treatment), it is worth noting that one can construct estimators for the $ATT(t)$ that are more precise than those in (5.2); see, e.g., Borusyak2023, Gardner2021, Harmon2024, Marcus_Santanna_2021, Wooldridge2021, and Chen_SantAnna_Xie_effientDiD_2024. In such cases, though, one would require parallel trends to hold pre-treatment periods, parallel pre-trends becomes directly testable (as the model is overidentified), and $\tau_{-k}$ could be used to assess its plausibility directly. Other types of overidentification tests could also be used in this case to assess the validity of the DiD model directly; see Marcus_Santanna_2021 and Chen_SantAnna_Xie_effientDiD_2024 for a discussion.
@@
The second lesson is that statistical precision shapes the usefulness of pre-trend estimates. The hypothesis tests for parallel pre-trends in Figure 3 are low-powered to detect practically important violations (Roth2022, and Freyaldenhoven_etal_2024). The $\tau_{-k}$ estimates fail to rule out flat pre-trends, but as we shall see below, they also fail to rule out large pre-trends that would indicate serious bias in the $\widehat{ATT}(t)$’s. They simply do not say very much. [^16] Another possibility is that very precisely estimated pre-trends are distinguishable from zero yet also rule out even small violations. The magnitude of the violations also matters: a precisely estimated but small violation of parallel trends in pre-treatment periods is “better” than an imprecisely potentially large estimated pre-trend that does not rule out zero. Roth2022 discusses how conditioning the analysis on these kinds of low-powered tests can exacerbate biases and should be interpreted with care.
@@
The third lesson is that researchers can make better use of pre-trend estimates by taking a stand on the size of plausible and/or problematic parallel trends violations. For example, Bilinski2018 propose selecting a value for differential trends that would fully explain the estimated treatment effects and then using it instead of zero as the null hypothesis when conducting statistical tests of the estimated pre-trends; see also Dette_Schumann_JBES_2024. Rambachan2023 develop inference methods for an approach that bounds the $ATT(t)$’s under assumptions about the maximum size of parallel trends violations based on pre-treatment periods (see also Manski2018). One can either select a magnitude using contextual knowledge or set it equal to a multiple of the largest period-to-period pre-trend estimate. One then constructs an identified set that contains $ATT(t)$ not under parallel trends but under the weaker assumption that parallel trends is not violated by more than this pre-determined maximum in each post-period. Importantly, Rambachan2023 also show how to use the estimated covariance matrix of the event study estimates to construct confidence intervals around the set. [^17] In cases where a measured variable accounts for the pre-trends, Freyaldenhoven2019 develop two-stage-least-squares estimators that recover $ATT(t)$ parameters by extrapolating the pre-period relationship between outcomes and the covariate into the post-period. Given the existence of these methods, which are theoretically grounded in how to consider violations of pre-trends, we caution producers (and consumers) of DiD work against the common practice of using only a simple “eye-test” for whether pre-trends differ substantially from zero.
@@
In our application, Rambachan2023’s method underscores how little information the pre-trend estimates convey. The largest one-period pre-trend in Figure 3 is between event-time -5 and -4, when outcomes fall by roughly four deaths more in the expansion group versus the non-expansion group. If we assume that parallel trend violations are no bigger than this, the identified set for $ATT(2014)$ is $-2.6\pm 4=[-6.6,1.4]$, and given the size of the pre-period standard errors, we obtain a robust confidence interval of $[-11.1,5.1]$, which spans implausibly large effects in both directions. Assuming smaller parallel trend violations would shrink this interval, but applying the method to subsequent event-times widens it. In general, Rambachan2023 provide a flexible method to use information about potential parallel trends violations drawn either from external knowledge or the $\tau_{-k}$ estimates, as well as the precision of the pre-trend estimates, to gauge (statistically) the robustness of the $\widehat{ATT}(t)$’s.
@@
When pre-trends suggest that Assumption PT-ES fails, a way forward is to assume that it holds only after conditioning on covariates and proceeding similarly to Section 4.4; we discuss this path in detail in Section 5.1.4. Alternatively, one can attempt to parametrically model the violations of parallel trends. Usually this is done by including unit-specific linear trends; see, e.g., Mora_Reggio_ER_2019, Wooldridge2021, Lee_Wooldridge_2023, and Freyaldenhoven_etal_2024. We note, however, that the practice of using unit-specific linear trends deviates from the standard DiD procedures: it relies on alternative identification assumptions involving an explicit parametric model for unit-specific trends. We also note that sensitivity analysis procedures that do not rely on such models are available—Rambachan2023 is one example—and we encourage practitioners to consider them.